OpenStudy (anonymous):

What is the area of a square with a perimeter of 14?

OpenStudy (anonymous):

12.25

OpenStudy (anonymous):

14 / 4 = 3.5 and then you do 3.5 * 3.5 = 12.25

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

To get perimeter you add all four sides. So in order to get the number for all four sides you divide by 4,which gives you 3.5 and since it is a triangle all 4 sides are equal. So now we have to look for the area of it. We have all the sides so this will be pretty easy. So to find area you multiply two sides by each other. So it doesn't matter which sides you pick because they are all the same since it is a square. So 3.5*3.5=12.25. So the area of the square is 12.25